OpenStudy (anonymous):

How do you simplify 9/3-square root of 2?

OpenStudy (anonymous):

1.58578643763 does this say anything

OpenStudy (secret66):

lol

OpenStudy (anonymous):

No, it has to be in a fraction..

OpenStudy (anonymous):

These are my choices. I don't want the answer, but could you explain to me how to get the answer?

OpenStudy (solomonzelman):

\[\frac{9}{3}-\sqrt{2}\]\[3-\sqrt{2}\]

OpenStudy (solomonzelman):

I think we misunderstood the question can you retype it? here is the latex for equation-editor use. \frac{ }{ } fill in the empty spaces in brackets (in equation editor)

OpenStudy (solomonzelman):

for square rot use sqrt{ }

OpenStudy (solomonzelman):

\[\frac{9}{3-\sqrt{2}}\color{blue} {\times \frac{3+\sqrt{2}}{3+\sqrt{2}}}\]

OpenStudy (solomonzelman):

\[\frac{27+9\sqrt{2}}{7}\]

OpenStudy (solomonzelman):

B

OpenStudy (anonymous):

Oh okay! That makes sense. Thank you very much!!